### v1

# references.bib Weak Ties Explain Open Source Innovation

Abstract
In a real-world social network, weak ties (reflecting low-intensity, infrequent interactions) act as bridges and connect people to different social circles, giving them access to diverse information and opportunities that are not available within one’s immediate, close-knit vicinity. Weak ties can be crucial for creativity and innovation, as it introduces new ideas and approaches that people can then combine in novel ways, leading to innovative solutions and creative breakthroughs. Do weak ties facilitate creativity in software in similar ways?
In this paper, we show that the answer is “yes.” Concretely, we study the correlation between developers’ knowledge acquisition through three distinct interaction networks on GitHub and the innovativeness of the projects they develop, across over 38,000 Python projects hosted on GitHub. Our findings suggest that the diversity of projects in which developers engage correlates positively with the innovativeness of their future project developments, whereas the volume of interactions exerts minimal influence. Notably, acquiring knowledge through weak interactions (e.g., starring) as opposed to strong ones (e.g., committing) emerges as a stronger predictor of future novelty.
I Introduction
In the late 1960s and early 1970s, Mark Granovetter, then a PhD student at Harvard University, studied a random sample of men living in a Boston suburb who had recently changed jobs, to understand how they learned about their new job opportunities. His 1973 paper [granovetter1973strength], which would go on to become the most cited work in the social sciences, had one key finding: that while a significant majority of people found their jobs through personal contacts rather than through formal channels like advertisements or employment agencies, these contacts were acquaintances (weak ties) rather than close friends or family (strong ties). Granovetter’s resulting theory on the “strength of weak ties” explains how weak ties are more effective in job searches—they act as bridges and connect individuals to different social circles, giving access to diverse information and opportunities that are not available within one’s immediate, close-knit network. The theory continues to accrue scholarly attention and large-scale confirmation in the digital age [rajkumar2022causal].
Beyond job searches, the theory had far-reaching implications. For example, weak ties impact, the diffusion of information, social mobility, community organization, and, the focus of this paper, creativity and innovation. The greater diversity of information enabled by weak ties can be crucial for creativity and innovation, as it introduces new ideas and approaches that might not be available within a close-knit group of strong ties, where information is more likely to be redundant. People can then combine these ideas in novel ways, leading to innovative solutions and creative breakthroughs [baer2010strength, rost2011strength, michelfelder2013and]. The famous “water-cooler effect,” that advocates of in-person work environments often quote, relies on the same mechanism—these casual encounters, which are typically with weak ties (work acquaintances rather than close colleagues), can lead to the exchange of diverse ideas and perspectives that spark creativity [brucks2022virtual].
Do weak ties facilitate creativity in software in similar ways? More generally, how are structural characteristics of the information networks of software creators associated with the innovativeness of the software they will create? For example, is the software more innovative when it is created by developers with access to more diverse information networks?
In this paper we answer these questions quantitatively, through a large-scale empirical study of over 38,000 open-source Python projects hosted on GitHub. For every project in our sample, we start by reconstructing three project-to-project networks based on interactions (i.e., commits, issues, and GitHub star events) of its core developers with other projects in the past, spanning ties of varying strength. Using these networks we then compute proxies for the amount and diversity of knowledge accessible to the core developers, before working on the focal project, via their weak and strong ties. Finally, building on a recent framework viewing software innovation as the novel recombination of software packages [fang2024novelty], we propose a novel measure of software innovativeness and estimate regression models to quantify the association between this measure and various network characteristics.
Our results indicate that software innovativeness increases when actors have access to a greater diversity of projects through the three networks we analyze, whereas the volume of projects they interacted with in the past does not correlate with project innovativeness. Moreover, we observe that the diversity of projects and knowledge accessed through low-effort interactions (weak ties), as captured for instance by “starring” repositories, has a greater impact on subsequent project innovativeness compared to interactions involving higher-effort engagements (strong ties) like issue comments and commits.
Overall, our study makes several key contributions, including a novel measure of software innovativeness, and an empirically validated operationalization of tie strength derived from actions available on GitHub. But most notably, we begin to fill an important knowledge gap. Despite growing realization of the importance of better understanding and enabling creativity and innovation in software engineering [jackson2022team, jackson2024creativity, inman2024developer], there is a dearth of research and empirical evidence in our community on the mechanisms through which creativity and innovation emerge. Our work is the first to translate established social science theory on the strength of weak ties to the software engineering domain, and to provide strong empirical evidence that similar innovation-enabling mechanisms apply also in (open-source) software. We discuss more implications for platform design and open source practice in Section V.
II Theoretical Framework and Hypotheses
In this section we develop our research hypotheses, drawing on established social science theories and a rich literature on open-source development and the GitHub platform.
II-A Innovation as Novel Recombination
Innovation is a critical aspect of science and technology development, marked by profound changes that can disrupt established competencies and lead to significant advancements [verhoeven2016measuring]. Open-source software is a manifestation of open innovation [west2006challenges, west2006patterns], by fostering a collaborative, transparent, and user-driven development environment that uses the collective intelligence and resources of a globally-distributed community. This approach accelerates the pace of innovation and leads to more robust, adaptable, and widely adopted software solutions that are increasingly being recognized as critical digital infrastructure that powers our society [eghbal2020working].
There have been various attempts to characterize innovation in software, with early work [couger1992measurement] recognizing the dimensions of novelty and utility and more recent work also considering distinctions in functionality, aesthetics, and other relevant features [beghelli2020novelty]. One prevalent discourse within the scholarly literature on scientific and technological innovation asserts that innovation often arises from preexisting components of knowledge [uzzi2013atypical, shi2023surprising], with the reconfiguration of knowledge serving as a predominant mechanism for generating novel and useful outcomes [schumpeter1939business]. In open source, this framing led to recently establishing an empirical link between the integration of atypical combinations of software packages, indicating novelty, and the higher subsequent popularity of the corresponding software projects, as a manifestation of utility [fang2024novelty].
How to facilitate the emergence of innovation remains unclear, though a common argument centers on facilitating creativity, as a precursor. Creativity and innovation are closely related but distinct concepts—creativity is about generating novel ideas, while innovation is about applying those ideas to create real value. It is expected that a software team that is more creative has access to richer knowledge (e.g., knowledge about the problem and technical approaches to solving it), and is better able to combine pieces of that knowledge to create an innovative solution [du2007role].
Of course, there are many facets to creativity, including as an individual personality trait [kandler2016nature]. In this paper we focus only on one. Specifically, we argue that creativity at the team level is a function of the knowledge networks of the team members and, while we cannot measure creativity directly, we can expect to see differences in the teams’ creative outputs (i.e., more novel software), that are associated with differences in the structure of those knowledge networks.
II-B Knowledge Acquisition through Social Learning
Historically, and especially with the advent of “social coding” platforms like GitHub, participation in open-source development has been rife with opportunities for social learning [crowston2005social], i.e., the process by which individuals acquire new knowledge, behaviors, skills, or attitudes through observation, imitation, and interaction with others. First, learning, including by observing what others are doing [yang2021differential], has been [lakhani2005hackers, krishnamurthy2006intrinsic] and remains [gerosa2021shifting] among the most important motivations for people to contribute to open source. In addition, modern code hosting platforms offer many opportunities for users to interact (e.g., collaboration on a shared codebase, issue discussions, code review. Through social media-like functionality (e.g., “following” [blincoe2016understanding]) and many available signals that offer a high degree of transparency [dabbish2012social] (e.g., repository badges [trockman2018adding]), the platforms also facilitate users quietly observing and being influenced by others’ behaviors.
These interactions with other individuals [von2003community] and the engagement with the artifacts they create [cervetti2009increasing] serve as channels for information dissemination and knowledge exchange. Indeed, prior studies have documented many such social learning effects, including choosing an open-source license [peng2013learning, singh2013networks], discovering and adopting emerging tools [singer2014software, lamba2020heard], and learning new software design principles and programming skills by reviewing code authored by others [nandigam2008learning, alami2019does].
Therefore, we expect that an increase in the volume of interactions among developers should correspond to a higher amount of knowledge transfer, thus better preparing developers for future innovations. Thus, we hypothesize that:
H1. The more interactions developers have with other developers and projects, the more innovative their projects are.
[…]
H2. The greater the informational diversity of developers’ past interactions, the more innovative their projects are.
II-C The Strength of Weak Ties
That in a social network ties can have varying strength, reflective of real-world factors like the duration of shared interactions, emotional depth, level of intimacy, and amount of reciprocal exchanges [granovetter1973strength], should be no surprise. Consequently, one can thus label the ties involving infrequent and less intimate interactions as “weak,” and those characterized by higher frequency of interaction and intimacy as “strong.”
Where Granovetter’s theory [granovetter1973strength] was groundbreaking, however, is in explaining that weak ties play a fundamentally different structural role in networks. As it turns out, strong ties tend to develop among individuals within the same community, while weak ties are more common between individuals from different communities. Here, a community is defined as a group of people who are predominantly interconnected with each other. This pattern arises because individuals within the same community have numerous opportunities for interaction and often share many mutual friends. This shared social context leads to increased similarity in behaviors, interests, and other characteristics, thereby enhancing intimacy among individuals [granovetter1973strength]. Furthermore, Heider’s Cognitive Balance Theory [munroe2007cognitive] proposes that when two individuals are linked only by weak ties yet share many common friends, it creates a “psychological strain;” and such situations, characterized by cognitive imbalance, are typically avoided by people. Therefore, weak ties between individuals sharing many mutual friends or belonging to the same community are unlikely to persist over time [granovetter1973strength].
The bridging role of weak ties between people of different communities offers great advantages for the diffusion of novel information. Within the confines of a single community, individuals tend to become more similar through frequent social interactions, leading to the consumption and circulation of similar types of information. Consequently, the information disseminated with the strong ties connecting individuals of a single community often contains less novel content, and at times may even be redundant, rendering it less valuable overall (the discussion of echo chambers on social media [kitchens2020understanding] is a prime example of this degradation of information quality). In contrast, information acquired through weak ties, or cross-community connections, tends to originate more from individuals with diverse backgrounds and knowledge bases. This information is often unavailable within one’s own community, making it more likely to be novel and of greater value [granovetter1973strength].
The advantage of weak ties in accessing novel and valuable information has been empirically evaluated in various contexts, including job seeking [rajkumar2022causal, granovetter1983strength], generation of innovative ideas [burt2004structural], and the diffusion of news feed content [bakshy2012role]. Similarly, we expect that the software-related knowledge transferred through weak ties is more valuable for the creation of innovative projects. Thus, we hypothesize that:
H3. The more the informational diversity of developers’ past interactions can be attributed to weak ties, the more innovative their projects are.
III Methods
To test our hypotheses, we reconstruct three historical project-to-project interaction networks (encoding possible knowledge flows across ties of varying strength) for a large sample of open-source Python projects hosted on GitHub, compute vector representations for all the nodes, and compute measures of tie strength and tie diversity based on these networks and representations. Separately, inspired by Fang et al.’s [fang2024novelty] operationalization of software innovativeness as novel recombination of software packages, we propose and compute a novel measure of project innovativeness that is better able to capture context and is also more scalable. Finally, we bring the two parts together, using regression modeling on this longitudinal data to test our hypotheses about factors associated with higher levels of innovativeness.
III-A Initial Data Collection
For our study, we augment existing data from the World of Code (WoC) dataset [ma2019world] with additional data mined using the GitHub API. WoC is arguably the most comprehensive record of open-source projects and their commit activities. Conveniently, WoC also contains pre-computed historical package manager dependency information for all projects using the most popular programming language / package manager combinations—we use this information when computing project innovativeness.
To keep our analysis tractable, we focus only on projects primarily written in Python, i.e., projects where more than 50% of the source code files are Python files. We chose Python due to its immense popularity in open source [srinath2017python] and its widespread usage across various domains by developers from diverse backgrounds [saabith2019python]. While not allowing us to directly generalize our findings, the considerable variability in the Python ecosystem across many characteristics should ensure points of overlap with many other ecosystems.
Because many GitHub repositories are intended as code dumps rather than active development projects and to account for more of the intricacies of GitHub data [kalliamvakou2016depth], we further apply a few common sense filters. That is, we exclude forks and repositories with fewer than ten commits in total. In addition, we further exclude repositories created after 2022, so that we can observe enough of the projects’ development to make inferences about their innovativeness.
For this remaining set of projects we query the GitHub API for historical information on all their commits, issues, and stars (the latter two are not available in WoC); for each such event we record its author and timestamp. We use this data for our network construction, detailed in Section III-B below.
As a final preprocessing step, we used typical heuristics to identify and filter out bot accounts [dey2020detecting, wessel2018power], which would distort our networks and subsequent measures, as they tend to have lots of activity. To this end, we searched for keywords in user logins, such as -bot or -robot, and labeled those users as bots after manual review. Additionally, we manually inspected the profile descriptions of the top 100 most active user accounts in our sample, ordered by the number of commits, to ensure the removal of highly active bot accounts that may have escaped the naming convention heuristics.
III-B Network Construction
We theorized in Section II above that 1) the interactions developers have with each other and with each other’s artifacts, i.e., the ties they form, create opportunities for knowledge transfer and facilitate social learning; and 2) “strong” and “weak” ties may play different roles and may have different effects. To operationalize these concepts, we first construct three interaction networks where every node is a project (i.e., GitHub repository), and the links between two nodes encode different interactions the developers of one project had with the other project; links are directed. We operationalize “developers” and “interactions” as follows.
III-B1 Developers—Core Developers
We consider only the actions of core developers in a focal project, identified as those contributors who authored at least five percent of all commits in the repository, and contributed a minimum of ten commits to the project. Conceptually, in open source the core developers wield the most influence over a project’s technical design decisions and development trajectory. Thus, we expect that the knowledge they (as opposed to more peripheral contributors) had access to prior to developing a focal project has the most influence on the design of that project. Pragmatically, focusing on core developers has the added advantage of keeping the analysis tractable, as popular projects may have tens of thousands of peripheral contributors.
Note that we use the benefit of hindsight here: i.e., essentially we label developers as “core” already at the start of a project, based on the amount of contributions they eventually made. Other alternatives [robillard2024communicating] that don’t rely on hindsight are possible, like considering the founder of a project. However, these have other limitations; e.g., there can be many influential developers that join a project later. Network-based operationalizations of “core” versus “periphery” have also been proposed [joblin2017classifying]. However, prior work [joblin2017classifying] demonstrated that count-based and network-based measures generally agree.
III-B2 Interactions—Commits, Issues, Stars
We consider three types of actions that one can make in a project: contributing changes to the codebase by making commits, posting or participating in an issue discussion, and starring a repository. Besides varying in effort (with commits typically requiring the most effort and star gazing the least of the three), conceptually, the three should also reflect varying levels of knowledge flow from the project to the author of the action. While commits can vary in size [alali2008s], a successful commit often requires a significant understanding of the project’s technical details [von2003community], indicating a path for considerable knowledge flow between the project and the commit author. Issues threads contain discussions on various project-related topics, like feature suggestions, bug reports, and user problems [merten2015requirements]. Engaging in issue discussions can reflect a deep understanding of the project, which can be further enhanced through feedback and comments on the same thread. However, on average, we expect that participating in issue discussions requires (and reflects) a less deep understanding of the project compared to making changes to its codebase. Finally, starring a project is often done as a sign of appreciation, a bookmarking attempt, or an intent for later use [borges2018s]. Starring a repository indicates at least some awareness of the project, but on average probably much more basic understanding than the other two.
Consequently, we construct three directed topological project-to-project networks, one for each of the three interaction types, to represent possible flow of knowledge across projects. We record an edge if and only if a core developer of project interacted with project within one year prior to their initial commit to project , as illustrated in Figure 2. The weight of each edge represents the number of core developers from project who engaged with project in a specific type of interaction. The one-year window models that recent interactions may be more relevant.
III-B3 Validation
To further validate that interaction via commits, issues, and stars is a good proxy for ties of decreasing strength, we empirically compare the transitivity of the three networks. First, consider the illustration in Figure 3: if node is connected to and , a triad is formed. In social network theory, if ’s connections to and are both strong ties, the triad is unstable and will likely evolve into a triangle where , , and are all connected. This evolution occurs because and have more opportunities to interact and increase their mutual familiarity due to their strong ties with . Thus, networks characterized by strong ties typically exhibit a greater number of triangles, as opposed to triads, compared to those with weak ties [granovetter1973strength]. Formally, this can be computed using the measurement of transitivity, defined as . A higher transitivity value (number of triangles relative to triads) indicates a network of stronger ties.
Similarly, we compare the transitivity values of our three networks, summarized in Table I.111Since network transitivity is usually defined for undirected networks, we first convert our three networks (which are all directed) to their undirected form. That is, there is an undirected edge between nodes and if there is either an edge from to or from to . In addition, the reported number of nodes excludes isolates (i.e., nodes without edges) from each network, and the number of edges is calculated in the undirected version of each network. Overall, we observe approximately an order of magnitude () difference in transitivity values between each pair of networks. Specifically, the commit network displays the highest levels of transitivity, followed by the issue network, while the star network exhibits the lowest level. Thus, these findings are consistent with our theoretical understanding of tie strength.
| Interaction | #Nodes | #Edges | Transitivity () |
| Commits | |||
| Issues | |||
| Stars |
III-C Network Measures
Next we describe the variables we computed to capture a project’s position in the different networks.
III-C1 Using Out-Degree Centrality to Measure the Amount of Interaction and Accessible Knowledge
For every project, we compute its out-degree centrality within each of the three networks (commits, issues, and stars), defined as the total sum of the weights of all its out-edges. These out-edges represent the number of actions initiated by the core developers of the project towards other projects in the past, and are a proxy for the amount of knowledge from other projects available to the core developers of the focal project; this knowledge may end up influencing the design of the focal project.
This step resulted in computing the three variables listed in Table IV under Degree Variables (original).
III-C2 Using A Node2Vec Model to Measure the Diversity of Accessible Knowledge
Next, we compute a proxy for the informational diversity of projects to which a focal project is connected within each interaction network. This diversity metric reflects the extent to which the focal project can access knowledge from a variety of sources. While conventional metrics like network constraint [burt1997note] and betweenness centrality [brandes2001faster] are commonly used to measure network diversity, they often do not scale well and can be computationally intensive, especially when dealing with large networks such as those in our study. Moreover, these metrics typically consider both the in-degree and out-degree edges of the focal node, while our study relies solely on the out-degree to represent the possible flow of knowledge from other projects to the focal one.
As a result, we use a Node2Vec model [grover2016node2vec], a classical approach from the machine learning domain, to generate vector representations (embeddings) for each node in the network based on its topological position. Initially, we generate sequences of random nodes from the graph using a random walk process: Starting from a random node in the network, we sample the next node by randomly selecting another node in the network that is adjacent to the current one. The likelihood of selecting each node is proportional to the edge weight between the candidate next node and the current node. This process continues until we have selected enough nodes to complete the walk (in our study, each walk comprises 20 nodes). We repeat this process to generate multiple walks (or sequences of nodes).
Subsequently, we use a skip-gram model [mikolov2013efficient], commonly used for generating embeddings for words in natural language, on the generated walks to obtain embeddings for each node. Nodes that frequently appear together in the same walk, representing sets of projects that are related in some informational sense, tend to have vector representations that are closely situated in the embedding space.
Next, for each project in the network, we examine all other projects that receive a directed edge from the focal project (denoted as set ). Our knowledge diversity index is the pairwise distance between any two projects in :
where and are projects in , and and are their vector representations from the Node2Vec model. The distance between and is the negation of the cosine similarity of the two vectors.
We calculate the diversity index for each node in each network. However, the index is undefined for projects with an out-degree centrality less than two. Therefore, such projects will be excluded from our regression analysis when considering diversity metrics as independent variables.
This step resulted in computing the three variables listed in Table IV under Diversity Variables (original).
III-C3 Validation & Exploratory Analysis
One can expect that the three degree variables and the three diversity variables are correlated to some extent, since they are computed in the same way (albeit on different networks). To prevent issues of multicollinearity in our subsequent regression models, which can both complicate the interpretation of the estimated coefficients and reduce the statistical power of our model [daoud2017multicollinearity], we first use Principal Component Analysis (PCA) [wold1987principal] to decompose the correlated variables into orthogonal principal components.
Since they are designed to capture different concepts, we run PCA separately for the out-degree centrality variables across the three networks (referred to as the degree variables) and the structural diversity measurements across the three networks (referred to as the diversity variables).222Recall that for the diversity variables, the dataset (hence also PCA) is restricted to projects with an out-degree of at least two in all three networks, since otherwise the variables are undefined. As standard, before running PCA we first log-transformed all input variables and scaled them to a mean of zero and a standard deviation of one.
The PCA resulted in three principal components for the degree variables and another three for the diversity variables, with the proportion of variance explained by each component listed in Table II. Inspecting the table, we observe that the first two PCs cumulatively explain over 80% of the variance in both groups, thus we decided to retain only two principal components each for the degree and diversity variables. That is, we will use these PCs (listed under Degree Variables (post-PCA) and Diversity Variables (post-PCA) in Table IV) instead of the original variables in our regression models below.
| Out-deg. centrality | Diversity index | |||||
| PC1 | PC2 | PC3 | PC1 | PC2 | PC3 | |
| Variance Explained | 0.64 | 0.22 | 0.14 | 0.51 | 0.29 | 0.19 |
| Cumulative Variance | 0.64 | 0.86 | 1.00 | 0.51 | 0.81 | 1.00 |
| Out-deg. centrality | Diversity index | |||||
| PC1 | PC2 | PC3 | PC1 | PC2 | PC3 | |
| 0.60 | 0.45 | 0.67 | 0.63 | 0.36 | 0.69 | |
| 0.61 | 0.28 | 0.74 | 0.64 | 0.24 | 0.72 | |
| 0.52 | 0.85 | 0.11 | 0.43 | 0.90 | 0.08 | |
To interpret the two PCs we turn to Table III, which lists the loadings of the principal components onto the original variables.333Principal components are linear combinations of the original variables, with the loadings indicating the contributing coefficients. For example, the first PC for the degree variables, , uses the coefficients in the first column in Table III: . Inspecting the table we make the following observations. First, the loadings onto the first principal component (PC1) are relatively consistent across networks for both the degree and the diversity variables. Thus, we interpret PC1 to represent the average degree (or diversity of ties) of the project across the three networks. For example, drawing on our theoretical framework (Section II), a project for which is high is expected to have many ties (on average, across the three networks), i.e., its core developers should have more sources to draw inspiration from. Similarly, the ties of a project for which is high are expected to span more of the knowledge embedding space we used to estimate tie diversity, i.e., its core developers could have access to more varied sources to draw inspiration from.
Second, the loadings onto the second principal component (PC2) exhibit a descending trend across networks, again similarly for both the degree and the diversity variables. Notably, the highest loading on the degree (or diversity) comes from the star network, followed by the issue network, while the lowest comes from the commit network. Thus, we interpret PC2 to represent the strength of network ties in the degree (or diversity of ties) metric, i.e., the strength of weak ties. A project for which is high is expected to get more of its connectivity through the star network. Analogously, the diversity of knowledge accessible to core developers in a project for which is high can be attributed relatively more to the star network (weak ties) compared to the commit or issue networks (stronger ties).
It may seem counterintuitive to reason about PC1 and PC2 jointly, especially in a regression modeling framework like we apply below. What does it mean to vary PC2 from low to high, for instance, while holding PC1 fixed? Can PC1 and PC2 even vary independently? They can in theory, because they are orthogonal by construction, but are the combinations of low-high values of PC1 and PC2 even observable in the real-world data? After all, regression is an interpolation mechanism. We illustrate this space of interpretations of PC1 and PC2 for the diversity variables with the artificial examples in Figure 4 (see our replication package for real-world examples corresponding to each quadrant). In the top left quadrant (PC1 and PC2 both high), the project core developers can tap into a diverse knowledge space via either strong or weak ties. In the top right (PC1 high and PC2 low), the overall diversity of information accessible is just as high, but it’s the strong ties that are the source of it. In the bottom left (PC1 low and PC2 high), the overall diversity of information available is lower, but whatever diversity there is, it can be attributed mostly to the weak ties. Finally, in the bottom right (PC1 and PC2 both low), weak ties contribute little to an otherwise low diversity of information.
Going back to our hypotheses (Section II), we expect the left quadrants, where weak ties are strong, to be most associated with innovativeness. As an anecdote, consider the example in Figure 1 on the first page. The GeoNotebook project is a Jupyter notebook-based environment for interactive visualization and analysis of geographic data. Interestingly, GeoNotebook’s core developers have previously starred (without otherwise contributing to) many GitHub projects spanning a variety of seemingly related topics, including a collection of free APIs, example data science notebooks, reproducible workflows for Jupyter notebooks, visualization tools, time series analysis tools, geographic data analysis tools, and infrastructure tools for interactive packet manipulation. One can expect that this diverse space of ideas provided some inspiration, knowingly or unknowingly, for the design and implementation of GeoNotebook. Our analysis below tests to what extent there is evidence supporting this mechanism at scale, beyond this anecdote.
| Outcome Variables | |
| Innovativeness | The extent to which a project imports software packages that were typically not imported together in the past, computed based on all the packages imported until the end of year 2021. |
| Degree Variables (original) | |
| The sum of edge weights from the focal project to the others in the commit network. | |
| The sum of edge weights from the focal project to the others in the issue network. | |
| The sum of edge weights from the focal project to the others in the star network. | |
| Degree Variables (post-PCA) | |
| (H) | The first principal component resulted from the PCA analysis over three original degree variables, representing average network degree across the three networks. |
| The second principal component resulted from the PCA analysis over three original degree variables, representing the degree from weak ties. | |
| Diversity Variables (original) | |
| The diversity of projects to which the focal project has a directed edge in the commit network. | |
| The diversity of projects to which the focal project has a directed edge in the issue network. | |
| The diversity of projects to which the focal project has a directed edge in the star network. | |
| Diversity Variables (post-PCA) | |
| (H) | The first principal component resulted from the PCA analysis over three original diversity variables, representing average diversity of connections across the three networks. |
| (H) | The second principal component resulted from the PCA analysis over three original diversity variables, representing the diversity from weak ties. |
| Controls Variables | |
| A fixed effect variable indicating the year in which the project received its first commit. | |
| Org_owned | A binary variable indicating whether the project was owned by an organizational account or not. |
| The total number of stars that the project owner received on all other repositories they owned before the creation of the focal project. | |
| The total number of core developers in the project, as computed before the end of year 2021. | |
| The number of software packages that the project imported before the end of year 2021. | |
III-D Project Innovativeness Measure
fang2024novelty introduced a measure of a software project’s novelty, or innovativeness, as a function of the combinations of packages the project imports. For example, while certain libraries, such as numpy and tensorflow, are frequently used together, others, like numpy and requests, are less commonly combined. The measure estimates how atypical combining two packages is, and by extension how atypical is the overall set of packages being used as dependencies in a project, by comparison to what could be expected by random chance. Projects that import more atypically paired packages are deemed more innovative as they reuse packages in nontraditional ways.
[…]
Next, we determine the atypicality of a combination of two packages by computing the negation of the cosine similarity of their embeddings. Finally, we define our measure of project innovativeness as the average pairwise atypicality of all pairs of packages imported by a project.444Recall that we can only compute the atypicality score for projects that import at least two packages. Other packages are excluded from the analysis.
[…]
III-E Regression Modeling Considerations
To test our three hypotheses, we use fixed-effects linear regression (one data point per project), where the response variable is our measure of innovativeness, and our main explanatory variables are the degree and diversity principal components discussed above: captures H (volume of interactions the core developers had with other developers and projects in the previous year), captures H (diversity of information accessible), and captures H (strength of weak ties). The remaining principal component, , does not directly map to one of our hypotheses, but can help distinguish whether it’s the volume of weak ties, or their strength, that explains more variance in our outcome measure.
In addition, we control for several important confounding variables, as described in Table IV: the repository owner’s social standing (), the project team size (), the complexity of the codebase in terms of number of packages imported (), whether the project is owned by an organizational account (Org_owned), and the project age (, modeled as a categorical variable). For model fit and diagnostics we followed standard practice, e.g., we first log-transformed the variables with skewed distributions to reduce heteroskedasticity [kaufman2013heteroskedasticity]; see replication package.
IV Results
Next we present the main results from the regression analysis formally testing our hypotheses, summarized in Table V, as well as a series of robustness checks.
| Model I | Model II | Model III | Model IV | |
| Variables of interest | ||||
| (H) | 0.001∗∗∗ | 0.001∗ | 0.002∗∗∗ | |
| (0.0002) | (0.001) | (0.001) | ||
| 0.00002 | 0.0001 | 0.005∗∗∗ | ||
| (0.0002) | (0.001) | (0.001) | ||
| (H) | 0.007∗∗∗ | 0.008∗∗∗ | ||
| (0.001) | (0.001) | |||
| (H) | 0.005∗∗∗ | 0.007∗∗∗ | ||
| (0.001) | (0.001) | |||
| Controls | ||||
| Org_owned | 0.027∗∗∗ | 0.017∗∗∗ | 0.018∗∗∗ | 0.018∗∗∗ |
| (0.001) | (0.002) | (0.002) | (0.002) | |
| (log) | 0.003∗∗∗ | 0.001 | 0.0004 | 0.001∗ |
| (0.0002) | (0.0004) | (0.0004) | (0.0004) | |
| (log) | 0.014∗∗∗ | 0.019∗∗∗ | 0.018∗∗∗ | 0.019∗∗∗ |
| (0.001) | (0.002) | (0.002) | (0.002) | |
| (log) | 0.040∗∗∗ | 0.023∗∗∗ | 0.023∗∗∗ | 0.024∗∗∗ |
| (0.0003) | (0.001) | (0.001) | (0.001) | |
| Fixed effect | ||||
| X | X | X | X | |
| Observations | 602,375 | 38,164 | 38,164 | 38,164 |
| ∗p0.05; ∗∗p0.01; ∗∗∗p0.001 | ||||
IV-A Main Regression Results
Model I (all projects) and Model II (projects with a minimum out-degree centrality of two across all three networks, such that the diversity variables are well defined) shed light on the impact of degree variables. Model I reveals a significant positive effect from average degree centrality, whereas Model II shows a negative effect. Nevertheless, effect sizes for both models are relatively small. For instance, in Model I, transitioning a project’s from the 25th percentile (-1.08) to the 75th percentile (0.54) results in a mere 0.002 increase in project innovativeness. Considering that the 25th percentile of a project’s innovativeness score is -0.33, and the 75th percentile is -0.12 in our sample, the change in project innovativeness due to variations in average degree centrality is practically negligible. Therefore, H is not supported.
In Model III, we observe a positive association between the average diversity () of ties originating from the focal project and project innovativeness. Moreover, this effect is notably stronger—nearly seven times greater—compared to that of average degree in Models I and II. Transitioning a project from the 25th percentile (-0.89) to the 75th percentile (0.86) in terms of corresponds to an increase of approximately 0.012 in project innovativeness, representing roughly a 4% change in the distribution (e.g., a shift from the median project in terms of novelty to the 54th percentile). We conclude that H is supported.
Additionally, we observe a significant effect for the variable. This suggests that, when controlling for the average diversity across the three networks, higher diversity in the “weak ties network” (e.g., star network), or stronger weak ties, corresponds to increased project innovativeness. In practical terms, the effect size for the variable is comparable, albeit somewhat smaller, to that of the variable above. We take this as evidence that H is supported.
When incorporating both degree and diversity variables in the same regression model, as in Model IV, the positive effects of and remain significant. This suggests a consistent effect stemming from tie diversity and strength of weak ties. Nevertheless, we also note a significant negative effect for and after adjusting for project diversity. These negative effects suggest that, while holding tie diversity constant, higher out-degree centrality and more ties in the “weak network” are associated with decreased project innovativeness. This observation could be attributed to the increased constraints within a local community associated with higher out-degree, as discussed by Burt [burt2000network]. Such constraints can lead to a reduced inclination for combining packages of diverse functions within the project, consequently contributing to the reduced novelty of the project. More research is needed to further test these effects.
IV-B Robustness Checks
In addition to the regressions above, we test how robust our conclusions about the effects of and are to differences in operational decisions.
IV-B1 Does The Time of Measurement Matter?
We segmented our dataset by project inception year and conducted regression analyses with the specification of Model III in Table V on projects initiated in each respective year. Figure 5(a) shows the resulting estimated coefficients for the two diversity variables. As illustrated in the graph, the positive relationship between the average diversity of knowledge access across the three networks (represented by the blue line) remains consistent over the years, with the average estimated coefficient being statistically significantly above zero in most years. The estimated coefficient for appears to also increase in recent years, though this difference is not statistically significant.
The positive relationship between diversity from weak ties and project innovativeness also appears generally persistent (the red line in Figure 5(a)). The estimated coefficient is above zero in nine out of the ten years observed, and the statistically insignificant coefficient (i.e., the lower bound of the 95% confidence interval is below zero) is likely attributable to the limited number of projects in each annual cohort (i.e., fewer than 4,000 projects in each year before 2016).
IV-B2 Does The Project Size Matter?
Similarly, we split projects into subsets based on the number of core developers and assessed the impact of the diversity variables in each cohort. As shown in Figure 5(b), the effect of remains consistent across projects of varying team sizes, confirming the robustness of our findings. It is noteworthy that the effect of becomes insignificant for projects with four or more core developers. This may be attributed to the limited number of project samples within these categories.
IV-B3 Does The Project Ownership Matter?
Lastly, we re-estimated the effects of the diversity variables separately for projects owned by organizational and individual accounts. Our analysis (Figure 5(c)) shows that both average diversity across networks and diversity from weak ties have significantly positive effects for both categories. Furthermore, the difference in effect size between organization-owned and individual-owned projects is not statistically significant.
In conclusion, the positive relationships we observed between average knowledge diversity and project innovativeness, as well as between diversity from weak ties and project innovativeness, are generally robust.
V Discussion
Next we discuss our results in the broader context of the literature and highlight budding directions for future research.
V-A Summary of Main Results
Established social science theory describes the mechanisms through which weak ties in a social network can be particularly “strong”—they often help bridge disjoint parts of the network, thus providing access to diverse and novel information, which in turn can be used towards better outcomes. Our study provides the first empirical link (that we know of) between this theoretical mechanism and a practical and important software engineering outcome—the emergence of innovation in a software project. Our hypotheses assert that the many user-to-user and user-to-artifact interactions possible on the GitHub platform, ranging from high-effort ones like making changes to a codebase to relatively trivial ones like starring a repository, result in network ties of varying strength, along which knowledge can flow. And, in particular, that weak ties are instrumental for this knowledge flow, as the GeoNotebook anecdote in the Introduction (Figure 1) would suggest.
Using a sophisticated methodology to reconstruct interaction networks, estimate tie strength, compute network informational diversity, and estimate project innovativeness, coupled with a robust statistical analysis, we found clear evidence supporting two of our three hypotheses: it’s not the amount (H), but rather both the diversity of information available to a focal project’s core developers (H), as well as the extent to which this information diversity comes from weak ties (H), that explain some of the variance in how novel (atypical) the combination of software packages eventually imported in the focal project are.
V-B Implications
One way to interpret our results is that lurking on the GitHub platform (where we see star gazing as lurking rather than actively contributing) has quantifiable benefits. If indeed people draw inspiration from things they’ve seen and apply those ideas to something they will build next,555Note, our study does not have sufficient power to make causal claims. platform designers should consider facilitating this process more. For example, GitHub has a way of highlighting “what the community is most excited about today” on its Trending page. It could be beneficial if such algorithms that determine what’s trending (allegedly666https://medium.com/@DiggerHQ/how-we-got-into-github-trending-c281f3b06df9 currently based on star count growth rate) also take into account the overall informational diversity of the highlighted repositories, or even better, the contribution to enhancing the informational diversity for a given user, i.e., if the recommendations were personalized.
In addition, research has shown that developers tend to join, or pay attention to, projects that are technically familiar to them [dey2021representation, fang2023matching], and automated project recommendation tools have been proposed, that recommend projects to developers with related technical background [zhang2017devrec]. Our results suggest that such approaches maximizing technical similarity might be counterproductive, at least in terms of enabling innovation.
Our results also raise questions about the GitHub culture of (implicitly) rewarding code contributions the most, since one’s commit history tends to be most highly visible in the platform. Much has been written about how other developers [marlow2013impression] and even recruiters [capiluppi2012assessing] use such signals. In contrast, we are finding evidence that well-informed but not necessarily highly active developers may also be experts at their craft, at least insofar as the novelty of their output.
It’s also worth thinking about what tracking and giving credit to ideas, not just code artifacts, might look like in this domain. In scientific research, authors are expected to acknowledge (and cite) prior work that influenced their thinking. In turn, these citation networks encode invaluable information about scientific progress and scientific collaboration [fortunato2018science]. What might such a system look like in software development? Anecdotally, it seems like developers already use many ad-hoc ways to credit ideas, e.g., including links in source code comments [hata20199] and even referring to research papers in repository README files [wattanakriengkrai2022github] or code [inokuchi2019academia]. But we still don’t understand how common the practice is, how complete the mentions are, and what the supply chains of ideas look like.
Questions around the value of having access to diverse information have also come up in research focused on demographic team composition, e.g., around gender [vasilescu2015gender] and race/ethnicity [shameer2023relationship]. This work posits similar underlying mechanisms (surface-level demographic variables act as a proxy for deeper-level, less observable differences in backgrounds, skills, approaches to solve problems, etc) to the one we measure here more directly. Now that such measurements are becoming possible, it would be interesting to further test the relationships between team-level surface attributes, informational or network diversity, and outcomes.
Finally, we can’t help but wonder how the proliferation of AI-based assistants will impact the innovativeness of the software being created. A large language model has certainly seen more diverse information in its training than any individual developer. But is it trained to expose that diversity, e.g., when generating code snippets automatically, or might AI-generated code end up looking more like a regression to the mean?
VI Conclusions
Much like in Granovetter’s jobs search study we started from in the Introduction, wherein most often people found their jobs through weak-tie acquaintances rather than strong-tie friends or family, we conclude that open-source developers may have found useful ideas to help them create innovative projects through passive GitHub interactions, like star gazing, more often than through active ones like committing to a repository.
VII Data Availability
Our replication package is available online777https://github.com/icsesubmission/replication_package_icse25_novelty and it includes aggregated data in CSV format and R scripts that reproduce the tables and figures in the paper.